An object is dropped from a helicopter which is moving horizontally at a constant velocity of 45m/s^1 180m above the ground. What is the time taken for the object to reach the ground?

1
Expert's answer
2022-02-23T08:30:37-0500

The time taken for the object to reach the ground does not depend on the initial horizontal speed and equals:


"t=\sqrt{\frac{2h}{g} } =\sqrt{\frac{2\cdot 180}{9.8} } =6.06\space s"